Calculate the pressure exerted by 0.45 mole of gas contained in a 120 L vessel at 25 degree Celsius.


1
Expert's answer
2021-01-26T03:24:19-0500

Given,

"n=0.45 , V=120\ L,T=25\degree C =(25 +273)K" "=298K" and "R=0.0821\ L\ atm \ K^{-1}mole^{-1}"

Using ideal gas equation ,

"PV=nRT" and substituting the values,

"P\times 120=0.45\times 0.0821\times 298"

"\implies P=0.09175\ atm"